# Fine-tuning Neural Network Quantum States

### 12 Mar 2024 v1

@@ Abstract
Recent progress in the design and optimization of Neural Network Quantum States (NNQS) have made them an effective method to investigate ground-state properties of quantum many-body systems. In contrast to the standard approach of training a separate NNQS from scratch at every point of the phase diagram, we demonstrate that the optimization at a highly expressive point of the phase diagram (i.e., close to a phase transition) yields interpretable features that can be reused to accurately describe a wide region across the transition. We demonstrate the feasibility of our approach on different systems in one and two dimensions by initially pretraining a NNQS at a given point of the phase diagram, followed by fine-tuning only the output layer for all other points. Notably, the computational cost of the fine-tuning step is very low compared to the pretraining stage. We argue that the reduced cost of this paradigm has significant potential to advance the exploration of condensed matter systems using NNQS, mirroring the success of fine-tuning in machine learning and natural language processing.
@@
In the last few years, neural networks are also increasingly used in condensed matter physics to approximate low-energy states of many-body quantum systems, both for spin and fermionic models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 9]. In this context, Neural Network Quantum States (NNQS) parametrize the amplitude of a variational state |$\ket{\Psi_{\theta}}$ expandend in a proper basis {|$\{\ket{\bm{\sigma}}\}$, mapping input physical configurations $\bm{\sigma}$ to complex numbers $\braket{\bm{\sigma}}{\Psi_{\theta}}=\Psi(\bm{\sigma};{\theta})$. Within the Variational Monte Carlo framework the parameters ${\theta}$ of the state are optimized to minimize the variational energy $E_{\theta}=\braket{\Psi_{\theta}}{\hat{H}}{\Psi_{\theta}}/\braket{\Psi_{\theta}}{\Psi_{\theta}}$ [19].
@@
Viteritti et al. [20] recently introduced a new parametrization of NNQS which explicitly leverages the feature learning perspective. In this viewpoint, a deep neural network is used to construct a map from the space of the physical configurations to abstract representations in a feature space, where the determination of the low-energy properties of the systems is simplified, and then a simple shallow network transform these representations into complex numbers. The resulting architecture achieved state-of-the-art performance on one of the most famous benchmark problems of the field [21].
In this Letter, we address a key conceptual question that this approach raises: given a system that exhibits a phase transition, do the representations learned to approximate the ground state near the transition point generalize to other points of the phase diagram? This question holds significance from both theoretical and practical perspectives. On the one hand, it has the potential to enhance the interpretability of Neural Network Quantum States. On the other hand, it offers a practical advantage by circumventing the need to optimize the wave function from scratch for each point in the phase diagram.
@@
The choice of pretraining the model near the phase transition is related to the fact that we expect the ground state to be very entangled close to critical points, thus spanning a large portion of the Hilbert space [22]. On the contrary, in trivial phases, where only a few configurations have non-zero amplitudes, the ability of pretrained networks to generalise away from these phases will be limited. We apply this procedure on finite clusters and measure physical properties (e.g., order parameters) of various systems exhibiting, in the thermodynamic limit, phase transitions of different nature. In all the cases, the features extracted during the pretraining stage, close to transition points, lead to excellent results after the fine-tuning in all the other points of the phase diagram. This is an important conceptual result, that reflects how neural networks can capture the essential quantum fluctuations in the vicinity of a phase transition.
@@
Conclusions. We showed that, for several physical systems exhibiting phase transitions, pretraining a Neural Network Quantum State near the transition point yields a set of features that can be fine-tuned to obtain accurate descriptions of the phase diagram. In addition, the analysis of the feature space facilitated the extraction of valuable insights into the structure of the wave function, enhancing the interpretability of NNQS. Furthermore, the fine-tuning process proves to be computationally more efficient, both in terms of time and memory, compared to traditional approaches in which the full training is required at each point of the phase diagram. We contend that, akin to the prevalent approach in machine learning [42], the employment of pretrained networks holds great potential for advancing the exploration of physical systems through NNQS. This aligns with the findings of a recent study by Ref. [43], demonstrating that, in electronic-structure problems, a single wave function can be employed to investigate multiple compounds and geometries. Exploring extensions of this approach to perform fine-tuning across different physical models stands as a crucial topic for future studies, as well as the development of techniques to automatically identify the most expressive pretraining point. Another possible application could focus on approximating the real-time dynamics [44, 45] by adjusting only the parameters of the shallow output network, thereby solving the problem within the feature space rather than the configuration space.

### 18 Mar 2024 v2

@@ Abstract
Recent progress in the design and optimization of Neural-Network Quantum States (NQS) have made them an effective method to investigate ground-state properties of quantum many-body systems. In contrast to the standard approach of training a separate NQS from scratch at every point of the phase diagram, we demonstrate that the optimization of a NQS at a highly expressive point of the phase diagram (i.e., close to a phase transition) yields interpretable features that can be reused to accurately describe a wide region across the transition. We demonstrate the feasibility of our approach on different systems in one and two dimensions by initially pretraining a NQS at a given point of the phase diagram, followed by fine-tuning only the output layer for all other points. Notably, the computational cost of the fine-tuning step is very low compared to the pretraining stage. We argue that the reduced cost of this paradigm has significant potential to advance the exploration of condensed matter systems using NQS, mirroring the success of fine-tuning in machine learning and natural language processing.
@@
In the last few years, neural networks are also increasingly used in condensed matter physics to approximate low-energy states of many-body quantum systems, both for spin and fermionic models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 9]. In this context, Neural-Network Quantum States (NQS) parametrize the amplitude of a variational state |$\ket{\Psi_{\theta}}$ expandend in a proper basis {|$\{\ket{\bm{\sigma}}\}$, mapping input physical configurations $\bm{\sigma}$ to complex numbers $\braket{\bm{\sigma}}{\Psi_{\theta}}=\Psi(\bm{\sigma};{\theta})$. Within the Variational Monte Carlo framework the parameters ${\theta}$ of the state are optimized to minimize the variational energy $E_{\theta}=\braket{\Psi_{\theta}}{\hat{H}}{\Psi_{\theta}}/\braket{\Psi_{\theta}}{\Psi_{\theta}}$ [19].
@@
Viteritti et al. [20] recently introduced a new parametrization of NQS which explicitly leverages the feature learning perspective. In this viewpoint, a deep neural network is used to construct a map from the space of the physical configurations to abstract representations in a feature space, where the determination of the low-energy properties of the systems is simplified, and then a simple shallow network transform these representations into complex numbers. The resulting architecture achieved state-of-the-art performance on one of the most famous benchmark problems of the field [21].
In this Letter, we address a key conceptual question that this approach raises: given a system that exhibits a phase transition, do the representations learned to approximate the ground state near the transition point generalize to other points of the phase diagram? This question holds significance from both theoretical and practical perspectives. On the one hand, it has the potential to enhance the interpretability of Neural-Network Quantum States. On the other hand, it offers a practical advantage by circumventing the need to optimize the wave function from scratch for each point in the phase diagram.
@@
The choice of pretraining the model near the phase transition is related to the fact that we expect the ground state to be very entangled close to critical points, thus spanning a large portion of the Hilbert space [22]. On the contrary, in trivial phases, where only a few configurations have non-zero amplitudes, the ability of pretrained networks to generalise away from these phases will be limited. We apply this procedure on finite clusters and measure physical properties (e.g., order parameters) of various systems exhibiting, in the thermodynamic limit, phase transitions of different nature. In all the cases, the features extracted during the pretraining stage, close to transition points, lead to excellent results after the fine-tuning in all the other points of the phase diagram [23]. This is an important conceptual result, that reflects how neural networks can capture the essential quantum fluctuations in the vicinity of a phase transition.
@@
Conclusions. We showed that, for several physical systems exhibiting phase transitions, pretraining a Neural Network Quantum State near the transition point yields a set of features that can be fine-tuned to obtain accurate descriptions of the phase diagram. In addition, the analysis of the feature space facilitated the extraction of valuable insights into the structure of the wave function, enhancing the interpretability of NQS. Furthermore, the fine-tuning process proves to be computationally more efficient, both in terms of time and memory, compared to traditional approaches in which the full training is required at each point of the phase diagram. We contend that, akin to the prevalent approach in machine learning [43], the employment of pretrained networks holds great potential for advancing the exploration of physical systems through NQS. This aligns with the findings of a recent study by Ref. [44], demonstrating that, in electronic-structure problems, a single wave function can be employed to investigate multiple compounds and geometries. A similar framework has also been proposed for decoding quantum error-correcting codes with generative modeling [45]. Exploring extensions of this approach to perform fine-tuning across different physical models stands as a crucial topic for future studies, as well as the development of techniques to automatically identify the most expressive pretraining point [46, 47, 48, 49]. Another possible application could focus on approximating the real-time dynamics [50, 51] by adjusting only the parameters of the shallow output network, thereby solving the problem within the feature space rather than the configuration space.